The Total number of electrons present in 48 g Mg^2+ are?
1
Expert's answer
2020-11-14T13:51:39-0500

24 g of Mg contains NA atoms.

"N_A = 6.023 \times 10^{23}"

1 atom of Mg contains 12 electrons.

Mg2+ ion contains (12-2) = 10 electrons.

2NA Mg2+ ions contain "10 \times 2N_A = 20N_A \;electrons"